### v1

# Complexity of Tensor Product Functions in Representing Antisymmetry

Abstract
Tensor product function (TPF) approximations have been widely adopted in solving high-dimensional problems, such as partial differential equations and eigenvalue problems, achieving desirable accuracy with computational overhead that scales linearly with problem dimensions. However, recent studies have underscored the extraordinarily high computational cost of TPFs on quantum many-body problems, even for systems with as few as three particles. A key distinction in these problems is the antisymmetry requirement on the unknown functions. In the present work, we rigorously establish that the minimum number of involved terms for a class of TPFs to be exactly antisymmetric increases exponentially fast with the problem dimension. This class encompasses both traditionally discretized TPFs and the recent ones parameterized by neural networks. Our proof exploits the link between the antisymmetric TPFs in this class and the corresponding antisymmetric tensors and focuses on the Canonical Polyadic rank of the latter. As a result, our findings uncover a fundamental incompatibility between antisymmetry and low-rank TPFs in high-dimensional contexts and offer new insights for further developments.
1 Introduction
High-dimensional partial differential equations (PDEs) and eigenvalue problems frequently arise in scientific and engineering applications. In solving these problems, traditional numerical methods, such as finite difference and finite element methods, suffer from the curse of dimensionality, in that both the storage and computational costs grow exponentially with the problem dimension. For example, discretizing an $N$-dimensional domain with merely two grid points per dimension can result in a dense tensor of size $2^{N}$, requiring over 1 ZB of memory for direct storage in double precision when $N\geq 70$.
To circumvent the curse of dimensionality, tensor product functions (TPFs) have been proposed to approximate the original high-dimensional ones [1, 5, 19]. For a function $f$ defined on a Cartesian product domain ×$\bigtimes_{j=1}^{N}\Omega_{j}$, with $\Omega_{j}\subseteq\mathbb{R}^{d}$ ($j=1,\ldots,N$), its TPF approximation is of the form
$f(\bm{r}_{1},\ldots,\bm{r}_{N})\approx\sum_{i=1}^{p}\psi_{i1}(\bm{r}_{1})\cdots\psi_{iN}(\bm{r}_{N}),\quad\forall~{}\bm{r}=(\bm{r}_{1},\ldots,\bm{r}_{N})\in\bigtimes_{j=1}^{N}\Omega_{j},$
where $p\in\mathbb{N}$ is called the separation rank in the literature [5, 6] (see Remark 2.2 for a rigorous definition) and $\psi_{ij}$ is a function defined on $\Omega_{j}$ ($i=1,\ldots,p$, $j=1,\ldots,N$). Combined with the traditional discretization methods for $\{\psi_{ij}\}$, the TPF approximation requires a complexity growing linearly with respect to $N$. The foundation of TPFs can be traced back to the work of Schmidt [47] on the case of $N=2$, $d=1$ with $\psi_{ij}\in\mathcal{L}^{2}(\Omega_{j})$, known as the Schmidt decomposition. Up to this point, this construction has been developed to handle higher-dimensional settings and function spaces with more favorable properties [4, 5, 6, 12, 17, 18, 20, 21, 40, 42].
In recent years, neural networks have gained significant attention for their powerful approximation capabilities (see [11] and the references therein for a comprehensive review) and provided mesh-free solutions for PDEs (see, e.g., [9, 13, 28, 35, 38, 46, 48, 49, 55, 57]). Particularly in the high-dimensional regime, Wang et al. [52] introduce the tensor neural networks (TNNs) by utilizing neural networks to construct $\psi_{ij}$ in the TPF approximation, effectively eliminating the need for high-dimensional integrals (see Section 2.1 for an illustration). TNN approximations have demonstrated high accuracy across various problems [27, 32, 36, 51, 52, 53, 54]; for instance, with a separation rank of $p\leq 30$, they achieve errors of $\mathcal{O}(10^{-7})$ for a 20,000-dimensional Schrödinger equation with coupled quantum harmonic oscillator potential [27].
However, recent studies have highlighted the limitations of TNN approximations in addressing the electronic Schrödinger equation in quantum mechanics [34]. This problem differs from the aforementioned applications of TNN approximations in the additional antisymmetry constraints on the unknown functions. Mathematically, an antisymmetric function $f$ defined on $\Omega^{N}$ ($\Omega\subseteq\mathbb{R}^{d}$) satisfies
$f(\bm{r}_{1},\dots,\bm{r}_{i},\dots,\bm{r}_{j},\dots,\bm{r}_{N})=-f(\bm{r}_{1},\dots,\bm{r}_{j},\dots,\bm{r}_{i},\dots,\bm{r}_{N}),\quad\forall~{}\bm{r}_{i}\in\Omega,~{}i\neq j.$ (1)
In the electronic Schrödinger equation, the electronic wave function describes the behavior of electrons, whose antisymmetry is grounded in the fundamental Pauli exclusion principle for fermions [44]. In [34], the authors approximate the electronic wave function with TNNs and optimize the neural network parameters through a variational principle. To achieve the desired accuracy even on systems with as few as three electrons, they have to employ TNNs with $p=50$ and undergo extraordinarily lengthy optimization processes.
The challenge can stem from the representation of antisymmetry. On the one hand, it is not difficult to observe that the TPFs are not inherently antisymmetric by definition, which is also evidenced by our numerical results on many-electron systems (see Figure 1).
Figure 1: Numerical comparison of the TNN approximations with and without explicit antisymmetrization on the one-dimensional systems of $\mathrm{Li}$ (left) and $\mathrm{HeH}^{+}$ (right).
These results also indicate that explicitly incorporating antisymmetry by antisymmetrization (see Eq. (6) for its definition) into TPF approximations can considerably improve both accuracy and computational efficiency. Detailed experimental settings are provided in Appendix A. On the other hand, existing works on representing antisymmetric wave functions usually rely on determinant-based constructions, such as Slater-type wave functions [14, 15, 29, 31, 45, 50, 58], pairwaise wave functions [22, 23, 43], Pfaffian wave functions [2, 3, 16], explicit antisymmetrization [37], and some implicit constructions [8, 56]. Within the framework of TPF approximation, these constructions correspond to separation ranks of $\mathcal{O}(N!)$. Together, these findings motivate us to investigate the complexity of TPFs in representing antisymmetric functions, a topic that remains unexplored in the literature but can offer valuable insights for tackling problems of interest.
Contributions. In this work, we examine a class of TPFs in which the functions $\{\psi_{ij}\}$ reside in finite-dimensional spaces. This class covers both TPFs discretized using traditional methods and the TNNs with any fixed network architecture. By linking the antisymmetric TPFs in this class to the associated antisymmetric tensors and analyzing the canonical polyadic rank of the latter, we establish an exponential lower bound of $\mathcal{O}(2^{N}/\sqrt{N})$ for the minimum number of terms required in the antisymmetric TPF representations. Our theoretical results reveal the fundamental limitations of TPFs for problems with antisymmetry constraints, including the electronic Schrödinger equation in quantum mechanics.
Organization. This paper is organized as follows: Section 2 introduces fundamental concepts related to TPFs, TNNs, and antisymmetric functions and tensors. Section 3 presents our main theoretical results on the antisymmetric TPFs and their application to the antisymmetric TNNs with fixed network architectures. Finally, Section 4 discusses the implications of these findings for relevant applications.
Notations. Throughout this paper, scalars, vectors, and tensors are denoted by lowercase letters (e.g., $x$), bold lowercase letters (e.g., $\bm{x}$), and uppercase bold letters (e.g., $\mathbf{X}$), respectively. Operators are represented by uppercase script letters (e.g., $\mathscr{A}$). We use subscripts to indicate components or blocks of vectors and matrices. For example, $x_{i}$ denotes the $i$-th component of $\bm{x}$, and $\bm{r}_{i}$ denotes the $i$-th block of $\bm{r}$. For an $N$-order tensor $\mathbf{X}$, the entry indexed by $(k_{1},\dots,k_{N})$ is written as $\mathbf{X}(k_{1},\dots,k_{N})$. The support of $\mathbf{X}$ is denoted by $\mathrm{supp}(\mathbf{X}):=\{(k_{1},\ldots,k_{N}):\mathbf{X}(k_{1},\ldots,k_{N})\neq 0\}$.
We use $\mathcal{B}(\Omega)$ for the Banach space of functions defined on the domain $\Omega$. The Cartesian products of sets are denoted by either exponents (e.g., $\Omega^{N}$) or multiplication symbols (e.g., $\Omega_{1}\times\Omega_{2}$ and ×$\bigtimes_{j=1}^{N}\Omega_{j}$). We use the notation $\otimes$ or $\bigotimes$ to represent tensor products of vectors (e.g., $\bm{x}_{1}\otimes\bm{x}_{2}$ and $\bigotimes_{j=1}^{N}\bm{x}_{j}$), vector spaces (e.g., $\mathbb{C}^{K_{1}}\otimes\mathbb{C}^{K_{2}}$ and $\bigotimes_{j=1}^{N}\mathbb{C}^{K_{j}}$), functions (e.g., $\psi_{1}\otimes\psi_{2}$ and $\bigotimes_{j=1}^{N}\psi_{j}$), and function spaces (e.g., $\mathcal{B}(\Omega_{1})\otimes\mathcal{B}(\Omega_{2})$ and $\bigotimes_{j=1}^{N}\mathcal{B}(\Omega_{j})$). Note that the tensor product of two functions is defined as $(\psi_{1}\otimes\psi_{2})(\bm{r}_{1},\bm{r}_{2}):=\psi_{1}(\bm{r}_{1})\psi_{2}(\bm{r}_{2})$ for any $\bm{r}_{1}$ and $\bm{r}_{2}$. The tensor product of function spaces can be obtained in an analogous way (see Definition 2.1). The notation $\binom{K}{N}$ represents the binomial coefficient, defined as $\binom{K}{N}=K!/(N!(K-N)!)$. We denote the permutation group over $\{1,\ldots,N\}$ as $\mathcal{S}_{N}$. The notation $\mathrm{sgn}(\pi)$ gives the sign of the permutation $\pi$.
2 Preliminaries
This section presents the concepts and tools necessary for our main results. We begin by detailing the structures of TPFs including TNNs, which provide the core frameworks for approximating high-dimensional functions. Following this, we introduce the mathematical properties of antisymmetric functions and tensors. For the broadest generality and compatibility with real applications, all functions in this work are assumed to be complex-valued and defined on real spaces.
2.1 Tensor Product Function and Tensor Neural Network
We begin by defining the TPF from the general Banach spaces of functions.
Definition 2.1 (TPF [19]).
Let $\mathcal{B}(\Omega_{j})$ be a Banach space of functions defined on $\Omega_{j}\subseteq\mathbb{R}^{d}$ ($j=1,\ldots,N$). The space of TPFs on $\{\mathcal{B}(\Omega_{j})\}_{j=1}^{N}$ is defined as
$\bigotimes_{j=1}^{N}\mathcal{B}(\Omega_{j}):=\left\{f\in\mathcal{B}(\bigtimes_{j=1}^{N}\Omega_{j}):f=\sum_{i=1}^{p}\bigotimes_{j=1}^{N}\psi_{ij},~{}\text{where}~{}p\in\mathbb{N},~{}\psi_{ij}\in\mathcal{B}(\Omega_{j}),~{}\forall~{}i,j\right\},$ (2)
and each function $f\in\bigotimes_{j=1}^{N}\mathcal{B}(\Omega_{j})$ is called a TPF (on $\{\mathcal{B}(\Omega_{j})\}_{j=1}^{N}$).
Remark 2.1.
The notation $\bigotimes_{j=1}^{N}\mathcal{B}(\Omega_{j})$ typically stands for the closure of the right-hand side of Eq. (2) [19], ensuring completeness. Since we focus on the functions with explicit tensor product structures, we omit the closure and, with slight abuse of notation, still denote the resulting space by $\bigotimes_{j=1}^{N}\mathcal{B}(\Omega_{j})$.
Unless stated, we assume that $\Omega_{1}=\cdots=\Omega_{N}=\Omega$, and we write $\bigotimes^{N}\mathcal{B}(\Omega):=\bigotimes_{j=1}^{N}\mathcal{B}(\Omega_{j})$. Since $\bigotimes^{N}\mathcal{B}(\Omega)$ is dense in $\mathcal{B}(\Omega^{N})$ for any compact $\Omega\subseteq\mathbb{R}^{d}$ [19], TPFs can in principle approximate any function in $\mathcal{B}(\Omega^{N})$ to arbitrary accuracy.
To facilitate our analysis, we define the rank of a TPF.
Definition 2.2 (TPF rank).
The TPF rank of a function $f\in\bigotimes^{N}\mathcal{B}(\Omega)$, denoted by $\mathrm{rank}_{\mathcal{B}(\Omega)}(f)$, is defined as
$\mathrm{rank}_{\mathcal{B}(\Omega)}(f):=\min\left\{p\in\mathbb{N}:f=\sum_{i=1}^{p}\bigotimes_{j=1}^{N}\psi_{ij},~{}\text{where}~{}\psi_{ij}\in\mathcal{B}(\Omega),~{}\forall~{}i,j\right\},$
or in other words, the minimum number of terms for the TPF representation of $f$.
Remark 2.2.
The definitions of TPF rank and separation rank [4, 5, 6, 17] differ fundamentally. The separation rank of a function is typically used in the context of approximation [4, 5, 6, 17]: Given a tolerance $\varepsilon>0$, the separation rank of a function $f$ is any $p\in\mathbb{N}$ such that there exist $\psi_{ij}\in\mathcal{B}(\Omega)$ ($i=1,\ldots,p$, $j=1,\ldots,N$) satisfying
$\left\|f-\sum_{i=1}^{p}\bigotimes_{j=1}^{N}\psi_{ij}\right\|_{\mathcal{B}(\Omega^{N})}\leq\varepsilon$
The optimal (or minimal) separation rank of $f$ then refers to its minimum separation rank for the given tolerance. In contrast, the TPF rank is defined specifically for TPFs, capturing the minimal number of terms required for an exact TPF representation. Since our goal is to analyze the complexity of TPFs in representing antisymmetry, where no approximation is involved, it is essential to define the TPF rank.
The TNN can be viewed as a special case of TPF with $d=1$ and $\{\psi_{ij}\}$ parameterized by neural networks [52]. Each input $r_{j}\in\Omega\subseteq\mathbb{R}$ is independently processed through a fully connected subnetwork that outputs a $p$-dimensional vector:
[…]
where $\bm{\theta}_{j}$ represents the parameters of the $j$-th subnetwork. Let $\bm{\theta}$ collect the parameters from $N$ subnetworks. The TNN function $f_{\bm{\theta}}$ is then defined as
$f_{\bm{\theta}}:=\sum_{i=1}^{p}\bigotimes_{j=1}^{N}\psi_{ij}(\cdot;\bm{\theta}_{j}),$ (3)
Figure 2 illustrates the architecture of TNN.
Figure 2: Architecture of TNN.
In the context of quantum mechanics, to approximate the electronic wave function of an isolated, non-relativistic, time-independent, and spinless $N$-electron system in $\Omega\subseteq\mathbb{R}^{d}$, the TNN wave function ansatz follows a similar structure. Each electronic coordinate is represented by a $d$-dimensional vector $\bm{r}_{j}=(r_{j1},\ldots,r_{jd})^{\top}\in\mathbb{R}^{d}$, with each one-dimensional coordinate processed by an independent subnetwork, resulting in a total of $dN$ independent subnetworks [34]. The TNN wave function ansatz for this system is
$f_{\bm{\theta}}(\bm{r}):=\sum_{i=1}^{p}\prod_{j=1}^{N}\left(\prod_{k=1}^{d}\psi_{ijk}(r_{jk};\bm{\theta}_{jk})\right),\quad\forall~{}\bm{r}\in\Omega^{N},$ (4)
where $\bm{\theta}$ collects the parameters of $dN$ subnetworks.
2.2 Antisymmetric Functions
Given a domain $\Omega\subseteq\mathbb{R}^{d}$, the antisymmetric function $f$ defined on $\Omega^{N}$ changes sign when any two inputs are exchanged, as defined in Eq. (1). Equivalently,
$f(\bm{r}_{1},\dots,\bm{r}_{N})=\text{sgn}(\pi)f(\bm{r}_{\pi(1)},\dots,\bm{r}_{\pi(N)}),\ \forall~{}\pi\in\mathcal{S}_{N},~{}\bm{r}\in\Omega^{N},$ (5)
where $\mathcal{S}_{N}$ is the permutation group on $\{1,\dots,N\}$, and $\text{sgn}(\pi)$ is the sign of permutation $\pi$. The set of all antisymmetric functions in $\bigotimes^{N}\mathcal{B}(\Omega)$ is denoted by
[…]
The antisymmetrizer (also known as the antisymmetrizing operator) [39] on $\Omega^{N}$, denoted by $\mathscr{A}$, is a linear operator that antisymmetrizes functions. For any $f$ defined on $\Omega^{N}$, the action of $\mathscr{A}$ on $f$ is defined as
$\mathscr{A}[f](\bm{r}_{1},\dots,\bm{r}_{N}):=\dfrac{1}{N!}\sum_{\pi\in\mathcal{S}_{N}}\text{sgn}(\pi)f(\bm{r}_{\pi(1)},\dots,\bm{r}_{\pi(N)}),\quad\forall~{}\bm{r}\in\Omega^{N}.$ (6)
For a function $f\in\bigotimes^{N}\mathcal{B}(\Omega)$, it belongs to $\mathcal{A}(\bigotimes^{N}\mathcal{B}(\Omega))$ if and only if $f=\mathscr{A}[f]$. When $\bigotimes^{N}\mathcal{B}(\Omega)$ is a Hilbert space, the antisymmetrizer $\mathscr{A}$ acts as the orthogonal projection operator onto $\mathcal{A}(\bigotimes^{N}\mathcal{B}(\Omega))$.
Remark 2.3.
Let us revisit the TPF in Eq. (2). When $p=1$, it reduces to the Hartree product [24], an early wave function ansatz in the quantum mechanics literature that lacks antisymmetry. When $p=N!$ with $\psi_{i1}=\text{sign}(\pi_{i})\psi_{\pi_{i}(1)}$ and $\psi_{ij}=\psi_{\pi_{i}(j)}$ for $1<j\leq N$, where $\pi_{i}\in\mathcal{S}_{N}$ for $i=1,\ldots,N!$, the TPF becomes the Slater determinant [15, 50], a well-known wave function ansatz or ingredient in various Hartree-Fock-based models for electronic structure calculations [25]. If antisymmetrized as defined in Eq. (6), the TPF becomes a sum of $p$ Slater determinants, whose TPF rank should satisfy $1<\mathrm{rank}_{\mathcal{B}(\Omega)}(\mathscr{A}[f])\leq pN!$.
2.3 Antisymmetric Tensors
The space of $N$-order $(K_{1},\dots,K_{N})$-dimensional tensors, denoted by $\bigotimes_{j=1}^{N}\mathbb{C}^{K_{j}}$, is defined as
$\bigotimes_{j=1}^{N}\mathbb{C}^{K_{j}}:=\mathbb{C}^{K_{1}}\otimes\cdots\otimes\mathbb{C}^{K_{N}}.$
In the special case where $K_{1}=\cdots=K_{N}=K$, we denote this space as $\bigotimes^{N}\mathbb{C}^{K}$. Since the TPF rank is highly relevant to the Canonical Polyadic (CP) rank of tensor (as will be shown later), we provide the formal definition of the latter as follows.
Definition 2.3 (CP Rank [26]).
The CP rank of a tensor $\mathbf{X}\in\bigotimes_{j=1}^{N}\mathbb{C}^{K_{j}}$, denoted by $\mathrm{rank}_{\mathrm{cp}}(\mathbf{X})$, is defined as
$\mathrm{rank}_{\mathrm{cp}}(\mathbf{X}):=\min\left\{p\in\mathbb{N}:\mathbf{X}=\sum_{i=1}^{p}\bm{x}_{i1}\otimes\cdots\otimes\bm{x}_{iN},\ \bm{x}_{ij}\in\mathbb{C}^{K_{j}},~{}1\leq i\leq p,~{}1\leq j\leq N\right\}.$
In analogy to the antisymmetric function, an antisymmetric tensor changes sign with the exchange of any two indices. Following this, we denote the closed subspace of $\bigotimes^{N}\mathbb{C}^{K}$ for all antisymmetric tensors as:
$\mathcal{A}\left(\bigotimes^{N}\mathbb{C}^{K}\right):=\left\{\mathbf{X}\in\bigotimes^{N}\mathbb{C}^{K}:\mathbf{X}(k_{1},\ldots,k_{N})=\text{sgn}(\pi)\mathbf{X}(k_{\pi(1)},\dots,k_{\pi(N)}),\forall~{}\pi\in\mathcal{S}_{N}\right\}.$
The unique orthogonal projection operator onto $\mathcal{A}(\bigotimes^{N}\mathbb{C}^{K})$, with a slight abuse of notation, is also denoted by $\mathscr{A}$: For any $N$-order tensor $\mathbf{X}$,
$\mathscr{A}[\mathbf{X}](k_{1},\ldots,k_{N}):=\dfrac{1}{N!}\sum_{\pi\in\mathcal{S}_{N}}\text{sgn}(\pi)\mathbf{X}(k_{\pi(1)},\dots,k_{\pi(N)}).$
If a tensor $\mathbf{X}$ is of a tensor product form $\mathbf{X}=\sum_{i=1}^{p}\bm{x}_{i1}\otimes\cdots\otimes\bm{x}_{iN}$, then its projection onto $\mathcal{A}(\bigotimes^{N}\mathbb{C}^{K})$ can be directly calculated by
$\mathscr{A}[\mathbf{X}]=\dfrac{1}{N!}\sum_{\pi\in\mathcal{S}_{N}}\sum_{i=1}^{p}\text{sgn}(\pi)\bm{x}_{i\pi(1)}\otimes\cdots\otimes\bm{x}_{i\pi(N)}.$
The following proposition gathers some key properties of antisymmetric tensors.
Proposition 2.1 ([19]).
Let $\mathbf{X}\in\mathcal{A}(\bigotimes^{N}\mathbb{C}^{K})$. Then:
  • If there exist $i\neq j$ such that $k_{i}=k_{j}$, then $\mathbf{X}(k_{1},\ldots,k_{N})=0$.
  • If $\mathcal{A}(\bigotimes^{N}\mathbb{C}^{K})\neq\{\mathbf{0}\}$, then $K\geq N$.
  • If $K\geq N$, then $\dim\mathcal{A}(\bigotimes^{N}\mathbb{C}^{K})=\displaystyle\binom{K}{N}$.
To write out the basis of $\mathcal{A}(\bigotimes^{N}\mathbb{C}^{K})$, we introduce the multi-index set
$\Lambda:=\left\{\bm{k}=(k_{1},\ldots,k_{N}):1\leq k_{1}<\cdots<k_{N}\leq K\right\}.$ (7)
Notice that $\left|\Lambda\right|=\binom{K}{N}$, and the entries of an antisymmetric tensor $\mathbf{X}$ indexed by the multi-indices in $\Lambda$ are independent. For each $\bm{k}=(k_{1},\ldots,k_{N})\in\Lambda$, we define the corresponding basis tensor $\mathbf{E}_{\bm{k}}\in\mathcal{A}(\bigotimes^{N}\mathbb{C}^{K})$ as
$\mathbf{E}_{\bm{k}}:=\sum_{\pi\in\mathcal{S}_{N}}\text{sgn}(\pi)\bm{e}_{k_{\pi(1)}}\otimes\cdots\otimes\bm{e}_{k_{\pi(N)}},$ (8)
where $\bm{e}_{j}\in\mathbb{C}^{K}$ is the $j$-th standard unit vector in $\mathbb{C}^{K}$. By the definition of $\mathbf{E}_{\bm{k}}$, we have
$\mathbf{E}_{\bm{k}}(\ell_{1},\dots,\ell_{N})=\begin{cases}1&\text{if }\{\ell_{1},\dots,\ell_{N}\}\text{ is an even permutation of }\{k_{1},\dots,k_{N}\},\\ -1&\text{if }\{\ell_{1},\dots,\ell_{N}\}\text{ is an odd permutation of }\{k_{1},\dots,k_{N}\},\\ 0&\text{otherwise}.\end{cases}$ (9)
Let $c_{\bm{k}}:=\mathbf{X}(k_{1},\ldots,k_{N})$ for any $\bm{k}\in\Lambda$. The expansion of $\mathbf{X}\in\mathcal{A}(\bigotimes^{N}\mathbb{C}^{K})$ in the basis $\{\mathbf{E}_{\bm{k}}\}_{\bm{k}\in\Lambda}$ is
$\mathbf{X}=\sum_{\bm{k}\in\Lambda}c_{\bm{k}}\mathbf{E}_{\bm{k}}.$ (10)
When $K=N$, the multi-index set $\Lambda$ contains only one element, and thus there is only one basis tensor, denoted as
$\mathbf{E}:=\sum_{\pi\in\mathcal{S}_{N}}\text{sgn}(\pi)\bm{e}_{\pi(1)}\otimes\cdots\otimes\bm{e}_{\pi(N)}.$ (11)
This tensor is also known as the determinant tensor [10].
3 Main results
We state and rigorously prove our main theoretical results in this section. In particular, we focus on the finite-dimensional $\bigotimes^{N}\mathcal{B}(\Omega)$ and establish an exponential lower bound on the TPF rank of any nonzero antisymmetric function therein. This result is further applied to antisymmetric TNNs.
3.1 Complexity of Antisymmetric Tensor Product Functions in Finite-Dimensional Spaces
We consider the situation $\mathcal{B}(\Omega)=\mathcal{F}_{K}(\Omega)$, where $\mathcal{F}_{K}(\Omega)$ is a $K$-dimensional space spanned by the linearly independent basis functions $\left\{\varphi_{k}\right\}_{k=1}^{K}$ defined on $\Omega$, i.e.,
$\mathcal{F}_{K}(\Omega):=\text{span}\left\{\varphi_{1},\dots,\varphi_{K}\right\}.$
Remark 3.1.
It should be remarked that the finite-dimensional setting described above aligns with common practice. Typically, the functions $\{\psi_{ij}\}$ in TPFs are first parameterized using a set of basis functions, followed by the implementation of a finite-dimensional discretized version [1]. Moreover, the results established in this finite-dimensional setting are also applicable to the TNNs parameterized by neural networks, as we will demonstrate in the next subsection.
Before giving the main results, we provide the following lemma on the linear independence of TPFs.
Lemma 3.1.
For any $1\leq k_{1},\ldots,k_{N}\leq K$, let the function $\Phi_{N,\bm{k}}$ ($\bm{k}:=(k_{1},\ldots,k_{N})$) be defined as $\Phi_{N,\bm{k}}:=\bigotimes_{j=1}^{N}\varphi_{k_{j}}$. Then $\{\Phi_{N,\bm{k}}\}_{\bm{k}}$ are linearly independent.
Proof.
We prove by mathematical induction on $N$. For the case of $N=1$, the conclusion holds trivially since $\left\{\varphi_{k}\right\}_{k=1}^{K}$ are already linearly independent. Now assume that the conclusion holds for the case of $N=m-1$, that is, the functions $\{\Phi_{m-1,\bm{k}}\}_{\bm{k}}$ are linearly independent. Consider the case of $N=m$.
Suppose that there exist constants $\{c_{k_{1},\dots,k_{m}}\}$ such that
$\sum_{k_{1}=1}^{K}\cdots\sum_{k_{m}=1}^{K}c_{k_{1},\dots,k_{m}}\bigotimes_{j=1}^{m}\varphi_{k_{j}}=0.$
By rearranging terms, we obtain
$\sum_{k_{1}=1}^{K}\cdots\sum_{k_{m-1}=1}^{K}\left(\bigotimes_{j=1}^{m-1}\varphi_{k_{j}}\right)\otimes\left(\sum_{k_{m}=1}^{K}c_{k_{1},\dots,k_{m}}\varphi_{k_{m}}\right)=0.$
By induction, we conclude that
[…]
Since $\left\{\varphi_{k}\right\}_{k=1}^{K}$ are linearly independent, it follows that $c_{k_{1},\dots,k_{m}}=0$ for any $1\leq k_{1},\ldots,k_{m}\leq K$, which implies that the conclusion also holds for the case of $N=m$. The proof is complete. ∎
In this finite-dimensional setting, analyzing the antisymmetric TPF reveals a relationship between its TPF rank and the CP rank of an antisymmetric tensor. One of our main results is stated as follows.
Theorem 3.1.
Assume that $K\geq N$ and $f\in\mathcal{A}(\bigotimes^{N}\mathcal{F}_{K}(\Omega))\backslash\left\{0\right\}$. Then the TPF rank of $f$ satisfies
$\mathrm{rank}_{\mathcal{F}_{K}(\Omega)}(f)\geq\min_{\mathbf{X}\in\mathcal{A}(\bigotimes^{N}\mathbb{C}^{K})\backslash\left\{\bm{0}\right\}}\mathrm{rank}_{\mathrm{cp}}(\mathbf{X}).$
Proof.
By definition, the TPF rank of $f\in\bigotimes^{N}\mathcal{F}_{K}(\Omega)$ can always be achieved. Let $p:=\mathrm{rank}_{\mathcal{F}_{K}(\Omega)}(f)$. Then there exist functions $\{\psi_{ij}\}\subseteq\mathcal{F}_{K}(\Omega)$ such that $f$ can be written in the form $f=\sum_{i=1}^{p}\bigotimes_{j=1}^{N}\psi_{ij}$. Expanding each $\psi_{ij}$ in terms of the basis functions in $\mathcal{F}_{K}(\Omega)$, we obtain
$\psi_{ij}=\sum_{k=1}^{K}c_{ijk}\varphi_{k},$
where $\{c_{ijk}\}\subseteq\mathbb{C}$ are the expansion coefficients. Thus, the TPF $f$ satisfies
$f(\bm{r})$ $=\sum_{i=1}^{p}\prod_{j=1}^{N}\sum_{k=1}^{K}c_{ijk}\varphi_{k}(\bm{r}_{j})$
$=\sum_{i=1}^{p}\sum_{1\leq k_{1},\dots,k_{N}\leq K}\prod_{j=1}^{N}c_{ijk_{j}}\varphi_{k_{j}}(\bm{r}_{j})$
$=\sum_{1\leq k_{1},\dots,k_{N}\leq K}\left(\sum_{i=1}^{p}\prod_{j=1}^{N}c_{ijk_{j}}\right)\varphi_{k_{1}}(\bm{r}_{1})\bm{\cdot}\cdots\bm{\cdot}\varphi_{k_{N}}(\bm{r}_{N})$ (12)
for any $\bm{r}\in\Omega^{N}$. On the other hand, applying the antisymmetrizer to $f$, we obtain
$\mathscr{A}[f](\bm{r})$ $=\dfrac{1}{N!}\sum_{\tau\in\mathcal{S}_{N}}\text{sgn}(\tau)\sum_{i=1}^{p}\prod_{j=1}^{N}\sum_{k=1}^{K}c_{ijk}\varphi_{k}(\bm{r}_{\tau(j)})$
$(\text{Let }\pi=\tau^{-1},\ell=\tau(j),\text{ and thus }j=\pi(\ell))$
$=\dfrac{1}{N!}\sum_{\pi\in\mathcal{S}_{N}}\text{sgn}(\pi)\sum_{i=1}^{p}\prod_{\ell=1}^{N}\sum_{k=1}^{K}c_{i\pi(\ell)k}\varphi_{k}(\bm{r}_{\ell})$
$=\dfrac{1}{N!}\sum_{\pi\in\mathcal{S}_{N}}\text{sgn}(\pi)\sum_{1\leq k_{1},\dots,k_{N}\leq K}\left(\sum_{i=1}^{p}\prod_{j=1}^{N}c_{i\pi(j)k_{j}}\right)\varphi_{k_{1}}(\bm{r}_{1})\cdots\varphi_{k_{N}}(\bm{r}_{N})$
$=\sum_{1\leq k_{1},\dots,k_{N}\leq K}\left(\dfrac{1}{N!}\sum_{\pi\in\mathcal{S}_{N}}\text{sgn}(\pi)\left(\sum_{i=1}^{p}\prod_{j=1}^{N}c_{i\pi(j)k_{j}}\right)\right)\varphi_{k_{1}}(\bm{r}_{1})\cdots\varphi_{k_{N}}(\bm{r}_{N})$
[…]
$=\sum_{1\leq k_{1},\dots,k_{N}\leq K}\left(\dfrac{1}{N!}\sum_{\pi\in\mathcal{S}_{N}}\text{sgn}(\pi)\left(\sum_{i=1}^{p}\prod_{j=1}^{N}c_{ijk_{\pi(j)}}\right)\right)\varphi_{k_{1}}(\bm{r}_{1})\cdots\varphi_{k_{N}}(\bm{r}_{N})$ (13)
for any $\bm{r}\in\Omega^{N}$. Since $f$ is antisymmetric, we have $f=\mathscr{A}[f]$. Noticing that $\left\{\varphi_{k}\right\}_{k=1}^{K}$ are linearly independent, we conclude after comparing Eqs. (12) and (13) and using Lemma 3.1 that the coefficients $\{c_{ijk}\}$ satisfy
$\begin{cases}\displaystyle\sum_{i=1}^{p}\prod_{j=1}^{N}c_{ijk_{j}}=\dfrac{1}{N!}\sum_{\pi\in\mathcal{S}_{N}}\text{sgn}(\pi)\left(\sum_{i=1}^{p}\prod_{j=1}^{N}c_{ijk_{\pi(j)}}\right),\ \forall~{}1\leq k_{1},\dots,k_{N}\leq K,\\ \displaystyle\exists\ k_{1},\dots,k_{N},\ \text{s. t. }\sum_{i=1}^{p}\prod_{j=1}^{N}c_{ijk_{j}}\neq 0.\end{cases}$ (14)
Define a tensor $\mathbf{X}\in\bigotimes^{N}\mathbb{C}^{K}$ by $\mathbf{X}(k_{1},\dots,k_{N}):=\sum_{i=1}^{p}\prod_{j=1}^{N}c_{ijk_{j}}$, and let $\bm{x}_{ij}=\left(c_{ij1},\dots,c_{ijK}\right)^{\top}\in\mathbb{C}^{K}$ for $1\leq i\leq p$ and $1\leq j\leq N$. Then
$\mathbf{X}=\sum_{i=1}^{p}\bm{x}_{i1}\otimes\cdots\otimes\bm{x}_{iN}.$
Thus, Eq. (14) implies that $\mathbf{X}\in\mathcal{A}(\bigotimes^{N}\mathbb{C}^{K})$ and $\mathbf{X}\neq 0$. By the definition of the CP rank, it follows that
$\mathrm{rank}_{\mathcal{F}_{K}(\Omega)}(f)=p\geq\min_{\mathbf{X}\in\mathcal{A}(\bigotimes^{N}\mathbb{C}^{K})\backslash\left\{\bm{0}\right\}}\mathrm{rank}_{\mathrm{cp}}(\mathbf{X}).$
which completes the proof. ∎
Remark 3.2.
The condition $K\geq N$ is necessary to avoid the trivial case. If $K<N$, by the basic property of antisymmetric tensors (see Proposition 2.1), it must hold that $f\equiv 0$.
This theorem establishes that the TPF rank of any $f\in\mathcal{A}(\bigotimes^{N}\mathcal{F}_{K}(\Omega))\backslash\left\{0\right\}$ is not smaller than the lowest CP rank of nonzero tensors in $\mathcal{A}(\bigotimes^{N}\mathbb{C}^{K})$. To estimate this lowest CP rank, we turn to analyze the CP ranks of the basis tensors $\{\mathbf{E}_{\bm{k}}\}_{\bm{k}\in\Lambda}$, as suggested by Eq. (10). For this purpose, we introduce an existing lower bound of $\mathrm{rank}_{\mathrm{cp}}(\mathbf{E})$, where $\mathbf{E}$ is the determinant tensor in Eq. (11).
Lemma 3.2 ([10]).
Let $\mathbf{E}\in\mathcal{A}(\bigotimes^{N}\mathbb{C}^{N})$ be the determinant tensor defined in Eq. (11). Then
$\binom{N}{\lfloor\frac{N}{2}\rfloor}\leq\mathrm{rank}_{\mathrm{cp}}(\mathbf{E})\leq N!\cdot\left(\dfrac{5}{6}\right)^{\lfloor\frac{N}{3}\rfloor}.$
The antisymmetric basis tensors $\{\mathbf{E}_{\bm{k}}\}_{\bm{k}\in\Lambda}$ defined in Eq. (8) generalize the determinant tensor $\mathbf{E}$ to higher dimensions. We now demonstrate that their CP ranks are identical.
Theorem 3.2.
Assume that $K\geq N$. Let $\Lambda$ be the multi-index set defined in Eq. (7). For any multi-index $\bm{k}\in\Lambda$,
[…]